# Error thresholds of toric codes with transversal logical gates

### v2

@@ Abstract
The threshold theorem promises a path to fault-tolerant quantum computation by suppressing logical errors, provided the physical error rate is below a critical threshold. While transversal gates offer an efficient method for implementing logical operations, they risk spreading errors and potentially lowering this threshold compared to a static quantum memory. Available threshold estimates for transversal circuits are empirically obtained and limited to specific, sub-optimal decoders. To establish rigorous bounds on the negative impact of error spreading by the transversal gates, we generalize the statistical mechanical (stat-mech) mapping from quantum memories to logical circuits. We establish a mapping for two toric code blocks that undergo a transversal CNOT (tCNOT) gate. Using this mapping, we quantify the impact of two independent error-spreading mechanisms: the spread of physical bit-flip errors and the spread of syndrome errors. In the former case, the stat-mech model is a 2D random Ashkin-Teller model. We use numerical simulation to show that the tCNOT gate reduces the optimal bit-flip error threshold to $p=0.080$, a $26\%$ decrease from the toric code memory threshold $p=0.109$. The case of syndrome error coexisting with bit-flip errors is mapped to a 3D random 4-body Ising model with a plane defect. There, we obtain a conservative estimate error threshold of $p=0.028$, implying an even more modest reduction due to the spread of the syndrome error compared to the memory threshold $p=0.033$. Our work establishes that an arbitrary transversal Clifford logical circuit can be mapped to a stat-mech model, and a rigorous threshold can be obtained correspondingly.
I Introduction
@@
The threshold theorem [1, 2, 3, 4] is a cornerstone of quantum error correction (QEC), promising a path toward fault-tolerant quantum computation (FTQC) [5, 6, 7]. The theorem states that for a quantum device with a physical error rate below a non-zero threshold, a QEC scheme can suppress the logical error rate to an arbitrarily low level. For instance, the toric code used as a quantum memory has a well-established bit-flip error threshold of approximately 10.9% for perfect syndrome measurements, or 3.3% when including syndrome errors [8, 9, 10]. Encouragingly, recent hardware breakthroughs have pushed various quantum platforms into this sub-threshold regime, enabling demonstrations where error correction improves the fidelity of logical memory [11, 12]. These advances set the stage for the implementation of fault-tolerant logical computations [13, 14, 15, 16, 17]. Two dominant schemes for this task are transversal gates [18, 19, 20] and lattice surgery [21], each having its own pros and cons. Specifically, compared to lattice surgery, transversal gates require fewer physical qubits and a shallower circuit to implement. However, physical errors propagate across code blocks via physical entangling operations under the transversal gate. It is to be expected, therefore, that the presence of transversal gates will lower the error threshold. Hence, rigorously establishing how transversal gates affect the error threshold has vast implications for the pursuit of fault-tolerant computation.
@@
Although Ref. [22] proved the existence of a non-zero error threshold of transversal logical circuit, the actual error threshold has only been observed empirically for specific, practical decoders [23, 24, 25, 26, 27, 28, 29, 30]. Therefore, a decoder-agnostic maximal error threshold is much needed for transversal logic circuits. For quantum memories, these tasks are accomplished through the statistical mechanical (stat-mech) mapping [8, 9], which maps the physical error rate to temperature and disorder strength, and the threshold to the critical point of order-disorder phase transition. The mapping leverages the correspondence between the large code distance limit and the thermodynamic limit of the statistical mechanics model to establish a rigorous threshold. However, to this date, stat-mech models in the presence of transversal gates are not known. In fact, rigorous derivation of error threshold using stat-mech mapping are only available for static quantum memories (see Table 1 for a list of existing stat-mech mappings for QEC codes).
In principle, the maximal error threshold of transversal logical circuits can also be crudely estimated from circuit simulation with a most likely error (MLE) decoder [24, 22], due to its near-optimal performance for toric code. However, since finding the MLE in general is an NP-hard problem for logical circuits, such an empirical approach is limited to small code distances. By deriving a mapping between transversal logical circuit under error and a classical stat-mech model, and simulating the model with finite size scaling to infer the critical point in the thermodynamic limit, we aim to establish a rigorous threshold in the infinite code distance limit (thermodynamic limit of stat-mech model).
@@
In this work, we show that the logical circuit applying a tCNOT gate to two toric code blocks with persistent bit-flip errors and perfect syndromes maps to a random Ashkin-Teller (AT) model. We perform classical Monte-Carlo simulations of the random AT model, which shows that the gate lowers the optimal bit-flip error threshold of the target code block to $p=0.080$, corresponding to a $26\%$ decrease from the memory threshold. With syndrome errors, we show that the logical circuit maps to a 3D random 4-body Ising model with a plane defect inserted in the time slice when the gate is applied. We then argue, using reported numerical results on 3D random 4-body Ising model, that the spreading of syndrome errors causes a more modest decrease of at most 14% in the target block’s threshold, compared to the memory threshold.
@@
This paper is organized as follows. In Sec. II, we review basics of stat-mech mapping for toric code quantum memory. In Sec. III, we introduce the tCNOT gate and spacetime detectors across the tCNOT gate. We then define the first noise models of our interest, namely the persistent bit-flip error and derive the stat-mech model of the tCNOT gate with perfect syndromes. We conduct Monte-Carlo simulation of the stat-mech model and obtain the corresponding error threshold in this case. In Sec. IV, we derive the stat-mech model for the second noise model, where the bit-flip error is combined with errors in syndrome extractions, and estimate the reduction of error threshold due to the tCNOT gate. Finally, in Sec. V, we discuss the stat-mech model of a general transversal logical circuit and its implication in fault-tolerant quantum computations.
@@
Logical ops? Stabilizer code Error model syndrome error? Stat-mech model
@@
Correlated error No RBIM with further neighbor term [32]
Single-qubit coherent error No Random complex coupling AT model [33, 34]
Both coherent and incoherent bit flip No
@@
CSS code + transversal Clifford Generic Pauli Yes See Appendix D
Table 1: Existing stat-mech mapping for 2D and 3D QEC codes. RBIM: random link Ising model. AT: Ashkin-Teller. RnnbIM ($n=3,4,6$): random $n$-body Ising model. 3D R4bIM is also known in the literature as random plaquette Ising gauge theory.
II Stat-mech mapping for error threshold of quantum memories
Determining the error threshold using stat-mech mapping consists of two steps. The first step is to map the probability distribution of syndromes to the partition function of a stat-mech model. The second step is to identify the thermodynamic quantity in the stat-mech model that corresponds to the logical error rate. Since our derivation closely follows the derivations of quantum memories, we first review the stat-mech mapping for quantum memories for completeness, following Refs. [8, 9, 32].
Figure 1: Stat-mech mapping of toric code under errors. (a) Sketch of trivial $C\in S$ and non-trivial cycle $C\in L$ on top of an error $E$ in one toric code block, which triggers a pair of syndromes denoted by the two red dots. (b) Stat-mech model of toric code under bit-flip errors. Here the physical qubits of the toric code are marked by circles, and the weight-4 Z stabilizers are defined on the shaded plaquettes. A trivial cycle $C\in S$ can be parameterized by a domain wall of Ising spins $\{\sigma\}$ (marked by the arrows) on a 2D square lattice that is rotated by $45^{\circ}$, marked by dashed lines. Two ends of a link $l$ on this lattice are denoted by $l_{1,2}$. (c) The 3D stat-mech model in the case of syndrome errors for a single code block. The Ising spins live at the links of thd 3D cubic lattice, marked by the arrows. Some of the 4-body interaction terms in the Hamiltonian in Eq. (II.2) are marked by grey stars. The dashed lines at $t\in\mathbb{Z}+\frac{1}{2}$ form a 2D square lattice where Ising spins live on the sites, which is the same 2D square lattice in (b).
II.1 Stat-mech mapping with perfect syndromes
We first consider the case with perfect syndromes, where every stabilizer measurement returns the actual physical value of the stabilizer. In this case, the error that occurs in the code state consists solely of physical Pauli errors, which will be denoted as a Pauli operator $E$. The error model is defined as the probability distribution $\mathrm{prob}(E)$. Physically, the Pauli error is a quantum channel that maps the density matrix of the system $\rho$ to $\sum_{E}\mathrm{prob}(E)E\rho E^{\dagger}$. In this case, the logical error rate, $P_{\text{logical}}$, can be computed as follows [8]:
@@ II.1 Stat-mech mapping with perfect syndromes
where $C$ denotes a “cycle” of Pauli operators that do not trigger any syndromes when the stabilizers are measured [^2] We call $C$ a cycle since in 2D topological codes such an operator is usually supported on closed cycles.. That is, it either belongs to the stabilizer group of the code state, denoted by $S$, or the set of non-trivial logical operator, denoted by $L$. In this way, Pauli errors $E$ and $E\cdot C$ will trigger exactly the same set of syndromes. In the first equality, the expression in square brackets is the probability of logical error given that a physical error $E$ occurs, which is computed via the second equality. Without the knowledge of the ground truth about $E$, the optimal decoder guesses an error $E\cdot C$ according to the conditional probability $\mathrm{prob}(E\cdot C|E)$. A logical error will occur whenever our guess differs from the actual physical error by any logical operator, see Fig. 1(a). Averaging over all possible errors yields the logical error rate in Eq. (II.1).
The central step of stat-mech mapping is to rewrite the summation of the cycles $C$ in Eq. (II.1) into a partition function of a stat-mech model with quenched disorder. For toric code under single qubit bit flip error, the cycle $C$ can be parameterized using a set of Ising spins $\{\sigma\}$ living on the lattice sites of a 2D square lattice, see Fig. 1(b). Topologically trivial cycle $C\in S$ is mapped to a topologically trivial domain wall of Ising spins. The summation of $C\in S$ now becomes the following partition function:
@@
The error threshold now has a clear physical picture: the critical point of the order-disorder phase transition of the stat-mech model. For low error rate, the stat-mech model has a low effective temperature (i.e. the couplings in $H(\{\sigma\}|E)$ are strong), and the stat-mech model is in the ordered phase. Consequently, the domain wall free energy $F_{DW}$ scales with the system size. Therefore, the logical error rate is exponentially suppressed with increasing code distance. In contrast, the high error rate is mapped to the disordered phase of the stat-mech model at high temperature, where the domain wall free energy $F_{DW}$ is finite. Therefore, $P_{\text{logical}}$ remains finite even in the large code distance limit. Based on numerical simulations in Ref. [40], the error threshold is $p_{\text{th}}=0.109$.
II.2 Stat-mech mapping with syndrome errors
We then review the stat-mech mapping for the case of syndrome errors. A common way to model this is to assume that every stabilizer measurement has the probability $q$ being the opposite of its true physical value. In this case, one generally needs multiple rounds of syndrome extractions to suppress logical errors. Therefore, the corresponding stat-mech model becomes three dimensional [9]. Again, the core in the stat-mech mapping procedure is to represent the summations over $C$ in Eq. (II.1) in partition functions. In the case of toric code memory under bit-flip noise with syndrome errors, every cycle $C$ can be parameterized by a set of Ising spins $\{\sigma\}$, which lives on the links of the 3D cubic lattice, see Fig. 1(c). The corresponding stat-mech model is the 3D random 4-body Ising model(R4bIM), also known as the random plaquette Ising gauge theory [8, 9, 10]:
@@ II.2 Stat-mech mapping with syndrome errors
In terms of the stat-mech model, the bit-flip error threshold of toric code with $O(d)$ rounds of syndrome extraction is mapped to the confinement transition of the R4bIM at $d\to\infty$ with increasing $p$ and $q$. At low error rates, the stat-mech model is in the deconfined (Higgs) phase, signaled by the perimeter-law decay of the Wilson loop order parameter $\langle W(\gamma)\rangle\equiv\langle\prod_{l\in\gamma}\sigma_{l}\rangle\sim e^{-A|\gamma|}$, where $\gamma$ is a closed loop in the cubic lattice, $|\gamma|$ is its length, and $A>0$ is some constant. Here, $\langle\dots\rangle$ represents both averaging over the partition function $Z_{3D}(E)$ and then averaging over the disorder $E$. At high error rates, the model enters the confined phase where the expectation value of the Wilson loop $W(\gamma)$ decays with the area $\mathfrak{S}_{\gamma}$ that the loop $\gamma$ encircles: $\langle W(\gamma)\rangle\sim e^{-B\mathfrak{S}_{\gamma}}$, where $B>0$ is some constant. The critical point of the confinement transition is around $p^{*}\approx 0.033$ in the case of $p=q$ [10].
III Toric codes under tCNOT gate with perfect syndromes
We now generalize the stat-mech mapping to logical circuits involving the tCNOT gate between two toric code blocks. Physically, the tCNOT gate is implemented via physical CNOT gates between every pair of physical qubits (see Fig. 2(a)). Since bit-flip operation on the control qubit does not commute with CNOT gate, the error will spread. In particular, we consider a specific bit-flip error model in this circuit that manifests the error spreading of bit-flip error by the tCNOT gate, and study its effect on the error threshold. The logical error rate in this case will still have the form of Eq. (II.1). However, we now need to determine the form of the cycle $C$ in logical circuits, in order to sum over all the cycles. The cycle $C$ is still defined as a set of errors, such that for any error $E$, $E$ and $E\cdot C$ trigger the same set of syndromes. The major difference here is that the notion of syndrome needs to be generalized to incorporate the propagation of errors by the transversal logical gates. Therefore, we first introduce the bit-flip error model of our interest, and then review error syndromes across a tCNOT gate.
III.1 Persistent bit-flip error model

### v3

# Rigorous estimation of error thresholds of transversal Clifford logical circuits

@@ Abstract
The threshold theorem promises a path to fault-tolerant quantum computation, provided the physical error rate is below a critical threshold. While transversal gates efficiently implement logical operations, they propagate errors and can lower this threshold relative to a static quantum memory. In this work, we generalize the statistical-mechanical (stat-mech) mapping from quantum memories to logical circuits with transversal gates, thereby enabling rigorous, decoder-independent thresholds for fault-tolerant logical computation. We first demonstrate the framework for two toric code blocks undergoing a transversal CNOT (tCNOT) gate, quantifying two independent error-spreading mechanisms. For persistent bit-flip errors with perfect syndromes, the stat-mech model is a 2D random Ashkin-Teller model. Monte Carlo simulation and finite-size scaling show that the tCNOT reduces the optimal bit-flip threshold to $p=0.080$, a $26\%$ decrease from the toric code memory threshold $p=0.109$. With syndrome errors included, the circuit maps to a 3D random 4-body Ising model with a plane defect, yielding a conservative estimate $p\geq 0.028$, a modest $15\%$ reduction from the memory threshold $p=0.033$. Beyond the tCNOT gate, we derive stat-mech models for all transversal Clifford gates of the toric code, including the fold-transversal Hadamard and $S$ gates, and generalize the framework to arbitrary CSS codes, proving that each transversal gate modifies the stat-mech model only locally in time. By reducing threshold analysis of fault-tolerant logical circuits to the study of classical spin models with local defects, our framework provides a systematic, decoder-independent benchmark for near-term fault-tolerant architectures.
1 Introduction
@@
The threshold theorem [1, 2, 3, 4] is a cornerstone of quantum error correction (QEC), promising a path toward fault-tolerant quantum computation (FTQC) [5, 6, 7]. The theorem states that for a quantum device with a physical error rate below a non-zero threshold, a QEC scheme can suppress the logical error rate to an arbitrarily low level. For instance, the toric code used as a quantum memory has a well-established bit-flip error threshold of approximately 10.9% for perfect syndrome measurements, or 3.3% when including syndrome errors [8, 9, 10]. Encouragingly, recent hardware breakthroughs have pushed various quantum platforms into this sub-threshold regime, enabling demonstrations where error correction improves the fidelity of logical memory [11, 12]. These advances set the stage for the implementation of fault-tolerant logical computations [13, 14, 15, 16, 17]. Two dominant schemes for this task are transversal gates [18, 19, 20] and lattice surgery [21], each having its own pros and cons. Specifically, compared with lattice surgery, transversal gates require fewer physical qubits and a shallower circuit. However, physical errors propagate across code blocks via transversal gate operations, thereby inevitably lowering the error threshold. Rigorously establishing how transversal gates affect the error threshold has thus become a pressing question for the pursuit of fault-tolerant computation.
@@
Although Ref. [22] proved the existence of a non-zero error threshold for transversal logical circuits, the actual threshold values have only been observed empirically for specific, practical decoders [23, 24, 25, 26, 27, 28, 29, 30]. Decoder-specific thresholds underestimate the fundamental limit: different decoders yield different values, and none provides the decoder-agnostic optimal threshold that bounds the performance of any decoding strategy. For quantum memories, the optimal threshold is established through the statistical mechanical (stat-mech) mapping [8, 9], a powerful analytical framework that maps the physical error rate to temperature and disorder strength of a classical spin model, and identifies the threshold with the critical point of an order-disorder phase transition. This mapping leverages the correspondence between the large-code-distance limit and the thermodynamic limit of the classical model, enabling well-controlled finite-size scaling (FSS) analysis to pinpoint thresholds with high precision. The stat-mech mapping has been successfully applied to a variety of topological codes and error models (see Table 1), in every case yielding insights inaccessible to decoder-based empirical studies.
Despite this considerable body of work on quantum memories, the stat-mech mapping has not yet been extended to logical circuits involving transversal gates. All existing stat-mech mappings listed in Table 1 are restricted to static quantum memories without logical operations. The fundamental obstacle is that transversal gates dynamically propagate errors between code blocks, modifying the structure of undetectable error cycles in a way that has no analog in the memory setting. In principle, the optimal threshold of a transversal logical circuit can be crudely estimated by circuit simulation with a most likely error (MLE) decoder [24, 22], due to its near-optimal performance for the toric code. However, since finding the MLE is in general NP-hard for logical circuits, such an empirical approach is limited to small code distances, precluding controlled extrapolation to the thermodynamic limit.
@@
In this work, we overcome this obstacle and generalize the stat-mech mapping to logical circuits with transversal Clifford gates. We show that the transversal gate modifies the classical spin model only locally in time, introducing a permutation defect of the Ising spins at the time slice where the gate is applied. This structural insight enables, for the first time in the context of logical circuits, rigorous, decoder-independent threshold analysis via FSS. In the main text, we demonstrate this framework for two toric code blocks undergoing a transversal CNOT (tCNOT) gate, which serves as a representative example for entangling transversal Clifford gates. We consider two complementary error models. First, for persistent bit-flip errors with perfect syndromes, we map the decoding problem to a 2D random Ashkin-Teller (AT) model, and use Monte Carlo simulations and FSS to determine that the tCNOT gate reduces the optimal bit-flip error threshold of the target code block to $p=0.080$, a $26\%$ decrease from the memory threshold. Second, for bit-flip errors coexisting with syndrome errors, the logical circuit maps to a 3D random 4-body Ising model with a plane defect; leveraging previously reported numerical results [10], we conservatively estimate a threshold of $p=0.028$, implying a more modest $15\%$ reduction from the memory threshold.
@@
This paper is organized as follows. In Sec. 2, we review the stat-mech mapping for toric code quantum memory. In Sec. 3, we introduce the tCNOT gate and derive the stat-mech model for persistent bit-flip errors with perfect syndromes, followed by Monte Carlo simulation to determine the corresponding error thresholds. In Sec. 4, we derive the stat-mech model with syndrome errors and estimate the threshold reduction. Finally, in Sec. 5, we discuss the implications of our stat-mech framework for fault-tolerant quantum computation.
Going beyond the tCNOT gate, in the Appendices we complete the stat-mech mapping for all transversal Clifford gates of the toric code, including the fold-transversal Hadamard and $S$ gates (Appendix C), and further generalize the framework to arbitrary Calderbank-Steane-Shor (CSS) codes with transversal Clifford gates using a binary matrix formalism (Appendix D). In the most general setting, we prove that each transversal gate locally modifies the stat-mech model at the time it is implemented, thereby establishing the scalability of our approach to entire logical circuits.
@@
Logical ops? Stabilizer code Error model Synd. error? Stat-mech model
@@
Correlated No RBIM + further neighbor [32]
Single-qubit coherent No Random complex coupling AT [33, 34]
Coherent + incoherent No
@@
CSS code + transv. Clifford Generic Pauli Yes See Appendix D
Table 1: Stat-mech mappings for QEC codes. All prior works (top section) apply exclusively to static quantum memories without logical operations. This work (bottom section) provides the first stat-mech mappings for logical circuits with transversal gates. RBIM: random bond Ising model. AT: Ashkin-Teller. RnnbIM ($n=3,4,6$): random $n$-body Ising model. 3D R4bIM is also known in the literature as the random plaquette Ising gauge theory.
2 Stat-mech mapping for error threshold of quantum memories
Determining the error threshold via a stat-mech mapping comprises two steps. The first step is to map the probability distribution of syndromes to the partition function of a stat-mech model. The second step is to identify the thermodynamic quantity in the stat-mech model that corresponds to the logical error rate. Since our derivation closely follows the derivations of quantum memories, we first review the stat-mech mapping for quantum memories for completeness, following Refs. [8, 9, 32].
Figure 1: Stat-mech mapping of toric code under errors. (a) Sketch of a trivial $C\in S$ and a non-trivial cycle $C\in L$ on top of an error $E$ in one toric code block, which triggers a pair of syndromes denoted by the two red dots. (b) Stat-mech model of toric code under bit-flip errors. Here, the physical qubits of the toric code are marked by circles, and the weight-4 Z stabilizers are defined on the shaded plaquettes. A trivial cycle $C\in S$ can be parameterized by a domain wall of Ising spins $\{\sigma\}$ (marked by the arrows) on a 2D square lattice that is rotated by $45^{\circ}$, marked by dashed lines. Two ends of a link $l$ on this lattice are denoted by $l_{1,2}$. (c) The 3D stat-mech model for syndrome errors in a single code block. The Ising spins reside on the links of the 3D cubic lattice, as indicated by the arrows. Some of the 4-body interaction terms in the Hamiltonian in Eq. (2.2) are marked by grey stars. The dashed lines at $t\in\mathbb{Z}+\frac{1}{2}$ form a 2D square lattice where Ising spins live on the sites, which is the same 2D square lattice in (b).
2.1 Stat-mech mapping with perfect syndromes
We first consider the case of perfect syndromes, in which every stabilizer measurement returns the actual physical value of the stabilizer. In this case, the error in the code state consists solely of physical Pauli errors, which we denote by the Pauli operator $E$. The error model is defined as the probability distribution $\mathrm{prob}(E)$. Physically, the Pauli error is a quantum channel that maps the density matrix of the system $\rho$ to $\sum_{E}\mathrm{prob}(E)E\rho E^{\dagger}$. In this case, the logical error rate, $P_{\text{logical}}$, can be computed as follows [8]:
@@ 2.1 Stat-mech mapping with perfect syndromes
where $C$ denotes a “cycle" of Pauli operators that do not trigger any syndromes when the stabilizers are measured [^2] We call $C$ a cycle since in 2D topological codes such an operator is usually supported on closed cycles.. That is, it either belongs to the stabilizer group of the code state, denoted by $S$, or the set of non-trivial logical operators, denoted by $L$. Thus, Pauli errors $E$ and $E\cdot C$ will trigger the same set of syndromes. In the first equality, the expression in square brackets is the probability of a logical error given that a physical error $E$ occurs, which is computed via the second equality. Without the knowledge of the ground truth about $E$, the optimal decoder guesses an error $E\cdot C$ according to the conditional probability $\mathrm{prob}(E\cdot C|E)$. A logical error will occur whenever our guess differs from the actual physical error by any logical operator, see Fig. 1(a). Averaging over all possible errors yields the logical error rate in Eq. (2.1).
The central step of stat-mech mapping is to rewrite the summation of the cycles $C$ in Eq. (2.1) into a partition function of a stat-mech model with quenched disorder. For the toric code under a single qubit bit flip error, the cycle $C$ can be parameterized using a set of Ising spins $\{\sigma\}$ living on the lattice sites of a 2D square lattice, see Fig. 1(b). A topologically trivial cycle $C\in S$ is mapped to a topologically trivial domain wall of Ising spins. The summation of $C\in S$ now becomes the following partition function:
@@
The error threshold now has a clear physical picture: the critical point of the order-disorder phase transition of the stat-mech model. For low error rate, the stat-mech model has a low effective temperature (i.e. the couplings in $H(\{\sigma\}|E)$ are strong), and the stat-mech model is in the ordered phase. Consequently, the domain wall free energy $F_{DW}$ scales with the system size. Therefore, the logical error rate is exponentially suppressed as the code distance increases. In contrast, the high error rate is mapped to the disordered phase of the stat-mech model at high temperature, where the domain wall free energy $F_{DW}$ is finite. Therefore, $P_{\text{logical}}$ remains finite even in the large code distance limit. Based on numerical simulations in Ref. [40], the error threshold is $p_{\text{th}}=0.109$.
2.2 Stat-mech mapping with syndrome errors
We then review the stat-mech mapping for syndrome errors. A common way to model this is to assume that every stabilizer measurement has the probability $q$ of being the opposite of its true physical value. In this case, one generally needs multiple rounds of syndrome extractions to suppress logical errors. Therefore, the corresponding stat-mech model becomes three-dimensional [9]. Again, the core in the stat-mech mapping procedure is to represent the summations over $C$ in Eq. (2.1) in partition functions. In the case of toric code memory under bit-flip noise with syndrome errors, every cycle $C$ can be parameterized by a set of Ising spins $\{\sigma\}$, which lives on the links of the 3D cubic lattice, see Fig. 1(c). The corresponding stat-mech model is the 3D random 4-body Ising model(R4bIM), also known as the random plaquette Ising gauge theory [8, 9, 10]:
@@ 2.2 Stat-mech mapping with syndrome errors
In terms of the stat-mech model, the bit-flip error threshold of the toric code with $O(d)$ rounds of syndrome extraction is mapped to the confinement transition of the R4bIM at $d\to\infty$ with increasing $p$ and $q$. At low error rates, the stat-mech model is in the deconfined (Higgs) phase, signaled by the perimeter-law decay of the Wilson loop order parameter $\langle W(\gamma)\rangle\equiv\langle\prod_{l\in\gamma}\sigma_{l}\rangle\sim e^{-A|\gamma|}$, where $\gamma$ is a closed loop in the cubic lattice, $|\gamma|$ is its length, and $A>0$ is some constant. Here, $\langle\dots\rangle$ represents both averaging over the partition function $Z_{3D}(E)$ and then averaging over the disorder $E$. At high error rates, the model enters the confined phase where the expectation value of the Wilson loop $W(\gamma)$ decays with the area $\mathfrak{S}_{\gamma}$ that the loop $\gamma$ encircles: $\langle W(\gamma)\rangle\sim e^{-B\mathfrak{S}_{\gamma}}$, where $B>0$ is some constant. The critical point of the confinement transition is around $p^{*}\approx 0.033$ in the case of $p=q$ [10].
3 Toric codes under tCNOT gate with perfect syndromes
We now generalize the stat-mech mapping to logical circuits involving the tCNOT gate between two toric code blocks. Physically, the tCNOT gate is implemented via physical CNOT gates between every pair of physical qubits (see Fig. 2(a)). Since the bit-flip operation on the control qubit does not commute with the CNOT gate, physical errors spread from the control to the target block. We consider a specific bit-flip error model that manifests this error-spreading mechanism, and study its effect on the error threshold. The logical error rate still takes the form of Eq. (2.1). However, we now need to determine the form of the cycle $C$ in the presence of logical gates, in order to sum over all cycles. The key difference from the memory case is that the notion of syndrome must be generalized to incorporate the dynamical propagation of errors through the transversal gate. We therefore first introduce the bit-flip error model of interest and then review error syndromes for a tCNOT gate.
3.1 Persistent bit-flip error model